### v1

@@ III-B NSR-ES
The second set of experiments was conducted with the quality-diversity NSR-ES algorithm. Again, a replication experiment was performed – our implementation of the algorithm was used to train the feed-forward model. This gave us a baseline for our further experiments with the transformers. And in the same manner as before in the previous subsection, we trained the Decision Transformer from scratch using NSR-ES, again with four times larger population. We then concluded with testing whether seeding the training with a pretrained transformer accelerates the training. Again, just as in the previous case, when training from pretrained models, no virtual batch normalization was used and values of learning rate and noise deviation hyperparameters were both reduced to $0.01$. The results of these experiments can be seen in Figures 2(e) and 2(f).
@@
To compare this algorithm with the previous one with respect to the final performance of agents trained by this algorithm in terms of distance traveled, we can take a look at Figure 3. We can see that the performance of the final agents based on a Decision Transformer this algorithm yields is comparable to the performance of such agents trained by the objective-based OpenAI-ES. The performance of NSR-ES seeded with a pretrained Decision Transformer is similar to OpenAI-ES – when the training is successful – but it is not so reliable and sometimes it fails to train the agent.
@@ IV Discussion
As for the NSR-ES, it proved more successful. Yes, it requires more computation than OpenAI-ES, but it yields comparable results with just a threefold increase of computation, while incorporating novelty, and thus being, in theory, capable of overcoming local optima.
Finally, the unfortunate pretraining. In our previous work [13], we noted that using the pretraining when utilizing OpenAI-ES faced serious challenges and proved to be pretty much useless. Still, we hoped that when used in the context of novelty-based training, it could insert a previous knowledge and accelerate the training. However, this did not happen and such attempts proved to be futile.
A possible solution, which might be further explored in future work, may be as follows. We start by training a smaller, weaker model using NS-ES or NSR-ES, saving the behavior archive from this training. Then, we create a new metapopulation of larger models by behavior cloning towards the original metapopulation and continue the training with this new metapopulation and the saved behavior archive. This ensures that there is no reinventing the wheel. In other words, this makes sure that the simple behaviors already tried by the simple models are not developed again. We hypothesize that this was the core reason for the nonfunctionality of pretraining in NS-ES and NSR-ES, and hence this could improve the performance significantly. Another benefit is that we could even keep and further use the virtual batch normalization data of the original metapopulation, as the inputs remain the same and we can use it already during the behavior cloning. This eliminates one of the problems of the utilization of pretrained models identified in our previous paper [13]. The only remaining problem is that the new metapopulation would not be so robust – being trained by a gradient algorithm or by a behavior cloning, respectively – and hence reduced values for hyperparameters influencing the speed of training would be required, which would dampen the progress of further training. Maybe this might be solved by gradually increasing the hyperparameters to their original values during the training. If this last bit could be resolved, we believe, this approach would outperform training the large models from scratch, and so it would help to solve more complex problems requiring more complex models for action selection.
@@ V Conclusion
We inspected the ability of novelty-based – either novelty search or quality-diversity – evolution strategies, in our case NS-ES and NSR-ES, to train larger and more complex models than the simple feed-forward ones that are standard across the reinforcement learning literature, like Decision Transformers. Although the novelty search algorithm would require much more computing power, the quality-diversity algorithm proved to be quite successful in training these bigger models.

### v2

@@ III-B NSR-ES
The second set of experiments was conducted with the quality-diversity NSR-ES algorithm. Again, a replication experiment was performed – our implementation of the algorithm was used to train the feed-forward model. This gave us a baseline for our further experiments with the transformers. And in the same manner as before in the previous subsection, we trained the Decision Transformer from scratch using NSR-ES, again with a four-times larger population. We then concluded by testing whether seeding the training with a pretrained transformer accelerates the training. Again, just as in the previous case, when training from pretrained models, no virtual batch normalization was used and values of learning rate and noise deviation hyperparameters were both reduced to $0.01$. The results of these experiments can be seen in Fig. 2(e) and 2(f).
@@
To compare this algorithm with the previous one with respect to the final performance of agents trained by this algorithm in terms of distance traveled, we can again take a look at Fig. 3. We can see that the performance of the final agents based on a Decision Transformer this algorithm yields is comparable to the performance of such agents trained by the objective-based OpenAI-ES. The performance of NSR-ES seeded with a pretrained Decision Transformer is similar to OpenAI-ES – when the training is successful – but it is not so reliable and sometimes it fails to train the agent.
@@ IV Discussion
As for the NSR-ES, it proved more successful. Although it does require more computation than OpenAI-ES, it yields comparable results with only a threefold increase in computation (thrice the iterations) while incorporating novelty and thus being theoretically capable of overcoming local optima.
Finally, the unfortunate pretraining. When experimenting with the objective-based OpenAI-ES [16], we noted that using the pretraining with the OpenAI-ES faced serious challenges and proved to be pretty much useless. Still, we hoped that when used in the context of novelty-based training, it could insert a previous knowledge and accelerate the training. However, this did not happen and such attempts proved to be futile.
A possible solution, which might be further explored in a future work, may be as follows: We start by training a smaller, weaker model using NS-ES or NSR-ES, saving the behavior archive from this training. Then, we create a new metapopulation of larger models by behavior cloning towards the original metapopulation and continue the training with this new metapopulation and the saved behavior archive. This ensures that there is no reinventing the wheel. In other words, this makes sure that the simple behaviors already tried by the simple models are not developed again. We hypothesize that this was the core reason for the nonfunctionality of pretraining in NS-ES and NSR-ES, and hence this could improve performance significantly. Another benefit is that we could even keep and further use the virtual batch normalization data of the original metapopulation, as the inputs remain the same and we can use it already during the behavior cloning. This eliminates one of the problems of the utilization of pretrained models identified during experiments with OpenAI-ES [16]. The only remaining problem is that the new metapopulation might probably not be so robust – being trained by a gradient algorithm, by a behavior cloning – and hence reduced values for hyperparameters influencing the speed of training would be required, which would dampen the progress of further training. Perhaps this might be solved by gradually increasing the hyperparameters to their original values during the training. If this last bit could be resolved, we believe, this approach would outperform training the large models from scratch, and so it would help to solve more complex problems requiring more complex models for action selection.
@@ V Conclusion
We inspected the ability of novelty-based evolution strategies, in our case NS-ES and NSR-ES – a novelty search and a quality-diversity algorithm, respectively – to train larger and more complex models than the simple feed-forward ones that are standard across the reinforcement learning literature, like Decision Transformers. Although the novelty search algorithm would require much more computing power, the quality-diversity algorithm proved to be quite successful in training these bigger models.